# VascX Models: Model Ensembles for Retinal Vascular Analysis from Color Fundus Images

### v1

@@ Results
0.5 Comparison with publicly available models
@@
Model ChaseDB DRHAGIS HRF RETA FIVES Leuven-Haifa RS
@@
Model HRF-AV RITE Les-AV Leuven-Haifa RS
@@ A V A V A V A V A V
Model ORIGA PAPILA IDRiD ADAM PALM REFUGE2 RS
@@
Model IDRID ADAM PALM REFUGE2 RS
@@
Table 2: Performance comparison of VascX segmentation models. NA is used when the model was trained on the corresponding dataset and it is not possible to benchmark on out-of-distribution data.
Table 2 shows the results of our comparison using Dice score (segmentation) and Euclidean distance (fovea localization) as evaluation metrics. In general, these results reveal higher performance from VascX across datasets and tasks.
Differences are smallest for the vessel segmentation task, where all models achieve consistently high Dice scores, suggesting that model performance in this task may be saturated for current architectures.
For A/V segmentation the VascX model achieved higher Dice scores across all datasets, with slightly higher performance for veins than for arteries, as has been observed in previous work. Importantly, we observed large improvements in Rotterdam Study and Leuven-Haifa datasets, the only ones evaluated across the three models (Figure 1). The differences was $\sim 0.1$ for the Rotterdam Study, the most diverse of the datasets; suggesting increased model robustness.
Regarding disc segmentation, the largest difference was in the Rotterdam Study, where Automorph achieved a score of only 0.220. It must be noted that some of the failure cases were due to Automorph’s bounds detection algorithm failing on RS CFIs. The disc detection was however also prone to fail for discs located close to or crossing the bounds of the CFI; and for images for which the disc border had low contrast. A large difference was also observed in PAPILA, which has a 30 degree field of view. The low performance of Automorph is likely partly from lack of 30-degree training samples when compared to VascX.
@@
Qualitatively we observed a marked improvement in segmentation performance from VascX models. Figure 3 shows example model outputs from Rotterdam Study and Leuven-Haifa inputs (the two independent datasets). We observed more consistent disc detection from VascX, especially in cases where the disc was not clearly differentiated. A/V segmentation, while not flawless, produced visibly less continuity and mis-classification mistakes than both LWNet and Automorph. Importantly, many of the vessel crossings were correctly resolved on both artery and vein masks.
0.6 Effect of image quality
To more deeply characterize the performance of the model, we made use of the CFI quality estimation model to classify Rotterdam Study images into Good, Usable and Bad / Unusable categories. We limited this analysis to the Rotterdam Study set due to the other dataset being smaller and/or highly uniform in terms of image quality and capture conditions. Figure 4(a) displays the distribution of Dice scores over image quality bins, as assessed by the model. Figure 4(b) shows the Dice score distributions for optic disc and macula-centered images. Once again vessel segmentation models achieved high consistency even for CFIs classified as bad quality. In A/V segmentation, bad images had lower performance. It stands out that the largest difference between VascX and the rest is in usable or intermediate quality images. Furthermore, it is clear that Automorph disc segmentation struggled with RS images, likely due to its more limited training set [40]. Performance was low for both disc and macula-centered images.
@@ 0.6 Effect of image quality
Figure 4: Box plots displaying the distribution of model performance (Dice score) over image quality and anatomic region (AR) bins/slices of the Rotterdam Study. For vessel segmentation, the performance of the models is stable across quality and anatomic region bins. For artery-vein segmentation, the biggest improvement is in usable images. For disc segmentation, the model achieves remarkably stable performance across image quality and AR; while the images pose a challenge for Automorph.
0.7 Feature accuracy
The goal of our models is to serve as the backbone for explainable feature extraction systems. Therefore, we compared retinal vascular features extracted from grader labels, Automorph segmentations, and VascX segmentations. Once again for VascX, we made use of segmentations of the test set for each fold to ensure out-of-distribution inference, and therefore the performance of our models might be under-represented. We made use of a feature extraction implementation used in previous work [27] for a set of 14 features including tortuosity and central retinal equivalents (CRE). All of these features make use of artery-vein segmentation, and some (CREs, temporal angles, vascular density) use disc segmentation and fovea location to define regions. Vessel segmentations were not used.
Table 3 shows the results, where it is clear that VascX produces features that are closer in magnitude and more highly correlated to the ground truth.
Grader $\mu$ Mean $\Delta$ Pearson r
Automorph VascX Automorph VascX
Temporal Angle - A 120.848 6.435 4.625 0.673 0.760
Temporal Angle - V 127.208 6.070 4.089 0.701 0.822
CRE - A 10.223 1.529 0.989 0.699 0.854
CRE - V 15.455 2.005 1.150 0.777 0.880
Vasc. Density - A 0.055 0.013 0.006 0.693 0.865
Vasc. Density - V 0.063 0.010 0.005 0.762 0.894
Caliber median - A 2.908 0.667 0.593 0.666 0.668
Caliber std - A 1.829 0.380 0.278 0.612 0.795
Caliber median - V 2.928 0.636 0.590 0.728 0.675
Caliber std - V 2.382 0.431 0.250 0.691 0.883
Tortuosity - A 1.087 0.007 0.006 0.733 0.772
Tortuosity - V 1.086 0.006 0.005 0.568 0.690
Bifurcations - A 44.607 22.930 20.337 0.687 0.706
Bifurcations - V 51.107 23.690 22.693 0.757 0.791
@@ 0.7 Feature accuracy
Table 3: Comparison of features extracted from ground truth annotations and from VascX and Automorph segmentations on the Rotterdam Study artery-vein segmentation set (300 CFIs). Grader $\mu$ is the mean of the feature; mean $\Delta$ is the mean absolute error between ground truth features and features extracted from model outputs. Pearson coefficients are correlations between ground truth and model outputs.
@@ Discussion
We presented new model ensembles for CFI vasculature analysis and benchmarked against publicly available systems. We developed the set of models in VascX with the objective of operating consistently across devices and achieving performance improvements on the diversity of images found on real cohort studies and clinical datasets. We augmented public datasets (which usually consist of carefully selected CFIs from one or a few devices) with Rotterdam Study CFIs collected throughout more than three decades from 11 different imaging devices and with diverse pathology. We made use of well-known architectures and strong augmentations to achieve solid performance across tasks.
Our results show that our models outperform those in previous systems in every metric we evaluated. Our models achieve better segmentation performance across image quality levels and are similarly well-performing on disc and macula-centered images. We observed the largest performance improvement on A/V segmentation on CFIs of intermediate quality, likely common in large cohorts and many clinical datasets. Importantly, our models also outperformed Automorph and LWNet models on the highest quality images. A large improvement was observed in disc segmentation performance when compared to Automorph, likely thanks to the greater image diversity on our augmented dataset. Finally, our results showed that better segmentation performance translates into better vascular features, in line with the results of Fhima et al. [10]. VascX features show lower absolute error and higher correlation with features extracted directly from grader segmentations.
@@
Further work is required in several areas. There is room for improvement in artery-vein segmentation, especially regarding the connectivity of the vessel trees. Despite the improvements, our model still produces A/V misclassifications and gaps. Enforcing the expected tree structure in the segmentation is a notorious open challenge in deep learning. Incorporating specific loss functions and post-processing steps may provide improvements. However, the Dice score is not a metric suitable for the evaluation of connectivity. The development and evaluation of metrics that capture connectivity is therefore relevant in this line of work.
Classification of image quality using a one-dimensional image-level label, while useful, is not granular enough for some applications. Artifacts or defocus, for example, may affect part of the image. A CFI that is unusable for vascular analysis may not be unusable for optic disc analysis, for example. Or a CFI may be usable for analysis of the main vascular arches but not the micro-vasculature.
We expect that VascX models will serve as the backbone for CFI vascular analysis, both in large cohort data and in smaller datasets. To this end, we expect that the ease of use of our Pytorch pipeline will aid researchers in efficiently obtaining segmentation and derived features. It is our objective that VascX catalyzes work toward understanding the statistical power of vascular analysis pipelines to expose signals in different kinds of data.
Furthermore, our vessel, disc, and fovea localization models may be useful beyond vascular where robust localization of these structures is required (eg. automatic ETDRS grid placement, or vessel-based quality estimation).
Data Sharing
With this publication, we make available our model weights, training, and inference pipeline code.
@@ Acknowledgments
We acknowledge the Eyened team, and in particular the Reading Center, for their role in developing the algorithms. We acknowledge the Sinergia consortium for conceiving and supporting this project. This work was funded by the Swiss National Science Foundation grant no. CRSII5209510.
@@ Appendix A Dataset details
Table 4: Complementary details of the public dataset used in model development.
@@
PALM [7] 1179** $2124\times 2056$, $1444\times 1444$ Zeiss Visucam 500, Canon CR-2
REFUGE2 [13] 2000 $1634\times 1634$ $2124\times 2056$ Zeiss Visucam 500, Canon CR-2, Topcon TRC-NW400, Kowa
RS (ours) 1225 see Fig. 1
@@
REFUGE2 [13] 2000 $1634\times 1634$ $2124\times 2056$ Zeiss Visucam 500, Canon CR-2, Topcon TRC-NW400, Kowa
RS (ours) 10908 see Fig. 1

### v2

@@ Results
Comparison with Publicly Available Models
@@
Model ChaseDB DRHAGIS HRF RETA FIVES Leuven-Haifa Rotterdam
@@
Model HRF-AV RITE Les-AV Leuven-Haifa Rotterdam
@@ A V A V A V A V A V
Model ORIGA PAPILA IDRiD ADAM PALM REFUGE2 Rotterdam
@@
Model IDRID ADAM PALM REFUGE2 Rotterdam
@@
Table 2: Performance comparison of VascX segmentation models. NA is used for a model - dataset pair when the model was trained on the dataset and therefore it was not possible to benchmark this model - dataset pair on out-of-distribution data.
Table 2 shows the results of our comparison using Dice score (segmentation) and Euclidean distance (fovea localization) as evaluation metrics, which reveal consistently higher performance from VascX across all datasets. Note that not all model - dataset pairs were evaluated due to the model being trained on the corresponding dataset entirely.
For A/V segmentation VascX achieved higher Dice scores across all datasets, with slightly higher performance for veins than for arteries, as has been observed in previous work [15]. Importantly, we observed large improvements in the Rotterdam and Leuven-Haifa datasets, the only ones evaluated across the three models (Figure 1). The differences was $\sim 0.1$ for the Rotterdam dataset, the most diverse of the datasets; suggesting increased model robustness.
Regarding disc segmentation, the largest difference was in the Rotterdam dataset, where Automorph achieved a score of only 0.220. Many of these failure cases were due to Automorph’s bounds detection algorithm failing on Rotterdam CFIs. The disc detection was however also prone to fail for discs located close to or crossing the bounds of the CFI; and for images for which the disc border had low contrast.
@@
Qualitatively we observed a marked improvement in segmentation performance from VascX models. Figure 4 shows example model outputs from Rotterdam and Leuven-Haifa inputs (the two independent datasets). We observed more consistent disc detection from VascX, especially in cases where the disc was not clearly differentiated. A/V segmentation, while not flawless, produced visibly less continuity and mis-classification mistakes than both LWNet and Automorph. Importantly, many of the vessel crossings were correctly resolved on both artery and vein masks.
Effect of Image Quality
To more deeply characterize the performance of the model, we made use of the CFI quality estimation model to classify Rotterdam images into Good, Usable and Bad / Unusable categories. We limited this analysis to the Rotterdam dataset set due to the other dataset being smaller and/or highly uniform in terms of image quality and capture conditions. Figure 5(a) displays the distribution of Dice scores over image quality bins, as assessed by the model. Figure 5(b) shows the Dice score distributions for optic disc and macula-centered images. Once again vessel segmentation models achieved high consistency even for CFIs classified as bad quality.
@@
Figure 5: Box plots displaying the distribution of model performance (Dice score) over image quality and anatomic region (AR) bins/slices of the Rotterdam dataset. For vessel segmentation, the performance of the models is stable across quality and anatomic region bins. For artery-vein segmentation, the biggest improvement is in usable images. For disc segmentation, the model achieves remarkably stable performance across image quality and AR; while the images pose a challenge for Automorph.
Feature Accuracy
The primary goal of our models is to serve as the backbone for explainable vascular feature/biomarker extraction systems. Therefore we quantified the quality of such features. We compared features extracted from model segmentations (VascX, Automorph (AM) and LittleWNet (LWN)) with reference features extracted from the annotations in our development set, which we considered ground truth. . We quantified the quality of model features by measuring mean absolute error (MAE) and correlations with the ground truth features. We used the Rotterdam set as main evaluation set for this comparison due to its diversity and size. Due to our 5-fold cross-validation setup for training, for VascX we made use of segmentations generated within each fold to have independence of the test set. Therefore the performance of our models might be under-estimated when compared to the full ensemble.
We made use of a vascular feature extraction to extract a set of features including tortuosity and central retinal equivalents (CRE) from segmentation masks. See Supplementary Section B.0.1 for feature implementation details. The features are calculated using artery-vein segmentation masks, and some (CREs, temporal angles, vascular density) use disc segmentation masks and fovea locations to define regions. Because LWNet does not include optic disc segmentation weights and Automorph struggled with the disc segmentation of Rotterdam images (Table 2(c)), we used the same ground truth disc masks in both pipelines. The fovea location was also the same across models. Therefore, this comparison evaluates the effect of different models for artery-vein segmentation only.
Table 3 shows the results of the comparison, where we measured reduced mean absolute error and higher correlations when using VascX segmentations, across most features. Of 24 vascular features evaluated, 16 showed a significant improvement in MAE when compared to Automorph, and 22 when compared to LWNet. VascX also had the highest correlations with ground truth features in all but one case.
We report results of the same comparison on the Leuven-Haifa dataset (240 images with $30\degree$ FOV) in Table 5. These results show similar behavior. Although less features improved over Automorph in terms of MAE, all correlations favoured VascX.
Grader $\mu$ MAE Pearson r
VascX AM LWN VascX AM LWN
Temporal Angle - A 122.36 5.56 6.39* 8.63$\dagger$ 0.69 0.66 0.57
Temporal Angle - V 126.35 5.05 5.91 8.29$\dagger$ 0.77 0.73 0.65
CRE - A 10.53 0.93 1.49$\dagger$ 3.10$\dagger$ 0.79 0.57 0.49
CRE - V 15.83 1.11 1.94$\dagger$ 3.79$\dagger$ 0.82 0.70 0.51
Vasc. Density - A 4.25 0.33 0.85$\dagger$ 0.76$\dagger$ 0.89 0.76 0.61
Vasc. Density - V 4.90 0.35 0.87$\dagger$ 1.12$\dagger$ 0.88 0.64 0.60
Vessel Caliber [med] - A 4.23 0.60 0.64 1.36$\dagger$ 0.63 0.48 0.37
Vessel Caliber [std] - A 2.26 0.27 0.39$\dagger$ 0.90$\dagger$ 0.70 0.59 0.49
Vessel Caliber [med] - V 4.35 0.83 .72* 1.38$\dagger$ 0.49 0.54 0.38
Vessel Caliber [std]- V 3.15 0.26 0.47$\dagger$ 0.79$\dagger$ 0.84 0.69 0.64
Tortuosity [med] - A 1.09 0.0058 0.006 0.0059 0.81 0.73 0.76
Tortuosity [med] - V 1.09 0.0048 0.0053 0.0059* 0.71 0.57 0.52
Curvature [med] - A 9.08 1.07 1.27* 1.34* 0.93 0.90 0.86
Curvature [med] - V 10.25 1.11 1.40$\dagger$ 1.44$\dagger$ 0.84 0.84 0.83
Inflection count [med] - A 1.46 0.48 0.61* 0.53 0.39 0.26 0.30
Inflection count [med] - V 1.57 0.45 0.50 0.60* 0.43 0.32 0.26
Tortuosity [LW] - A 1.12 0.017 0.019* 0.038$\dagger$ 0.52 0.47 0.11
Tortuosity [LW] - V 1.11 0.012 0.014* 0.033$\dagger$ 0.75 0.62 0.12
Bif. Angles [mean] - A 82.99 5.79 6.70* 9.51$\dagger$ 0.48 0.41 0.27
Bif. Angles [med] - A 81.83 5.19 6.15 8.97$\dagger$ 0.57 0.45 0.22
Bif. Angles [mean] - V 84.16 4.47 5.53* 7.71$\dagger$ 0.75 0.59 0.54
Bif. Angles [med] - V 83.20 3.75 5.02* 5.96$\dagger$ 0.74 0.52 0.56
Num. Bifurcations - A 27.32 10.34 10.97 12.59$\dagger$ 0.81 0.77 0.69
Num. Bifurcations - V 31.61 10.33 12.06$\dagger$ 13.48$\dagger$ 0.89 0.76 0.76
@@
Table 3: Feature quality comparison on the Rotterdam sets (215 CFIs). Results of comparing MAE of features extracted using different models for artery-vein segmentation: VascX, Automorph (AM) and LittleWNet (LWN). Features extracted from ground truth segmentations were used as the reference for MAE. The feature implementation was the same across models. For vessel or bifurcation features, the aggregation function (applied on features from the whole image) is indicated in brackes: [med]: median, [std]: standard deviation, [mean]: mean value, [LW]: length-weighted; vessels were weighted by their length. Grader $\mu$ is the mean of feature values; mean MAE is the mean absolute error between ground truth features and features extracted from model outputs. Pearson coefficients are correlations between ground truth and model outputs. *: Significant difference in MAE ($p<0.05$) between the model and VascX. $\dagger$: $p<0.001$
@@ Discussion
We presented new model ensembles for CFI vasculature analysis and benchmarked against publicly available systems. With the objective of developing models that operate robustly across devices and achieve performance improvements on the different imaging conditions found on real cohort studies and clinical datasets, we focused on the diversity of our development set. We augmented more than 15 public datasets (each of which usually consist of carefully selected CFIs from one or a few devices) with CFIs collected in The Netherlands throughout more than three decades from 11 different imaging devices and with diverse pathology. Together, this resulted in a comprehensive set of images including a wider range of patient characteristics (origin, age, pathologies) and capture characteristics (device, anatomic region, field of view). For model training, we made use of well-known architectures, a new more robust pre-processing algorithm, and data augmentation to achieve solid performance across tasks.
Our main evaluation using Dice scores revealed consistently better performance from VascX when compared to existing publicly available models - Automorph and LWNET - across both public datasets and the Rotterdam sets. We observed an important improvement in A/V segmentation performance, crucial in vascular analysis pipelines. It stands out that the largest difference between VascX and the rest is in usable or intermediate quality images 5, common in large cohort studies and clinical data. A large improvement was observed in disc segmentation performance when compared to Automorph, which struggled with Rotterdam images likely due to its more limited training set [57]. Its performance was low for both disc and macula-centered images and across quality levels (5). A large difference was also observed in PAPILA 2, which has a 30 degree field of view not present in Automorph’s training set. The differences in vessel segmentation performance were smaller between models, suggesting that performance in this task may already be saturated.
Our fovea localization model had mean L2 errors between 10.68 (REFUGE2) and 26.22 (PALM) on images of $1024\times 1024px$. This model is enough to provide an approximate localization of the fovea center, useful for the automatic placement of retinal regions (eg. ETDRS grid) for localized measurements.
Importantly, our results showed that better segmentation performance translates into more accurate vascular features, in line with the results of Fhima et al. [15]. Features from VascX model segmentations show lower absolute error and higher correlation with features extracted directly from grader segmentations; when compared to features extracted from Automorph segmentations.
@@
Further work is required in several areas. There is room for improvement in artery-vein segmentation, especially regarding the connectivity of the vessel trees. Despite the improvements, our model still produces A/V mis-classifications and gaps. Enforcing the expected tree structure in the segmentation is a notorious open challenge in deep learning. Incorporating specific loss functions [44] and post-processing steps [5] designed for connectivity may help address this issue.
The further development of retinal analysis pipelines and the development and evaluation of vascular features with clinical significance also holds promise. The feature computation stage involves significant nuance. Some features such as tortuosity, for example, may be particularly sensitive to discontinuities in the segmentation. The results in Table 3 support that some features are more robust than others to potential mistakes in the segmentations. Deeper evaluations of the robustness and predictive power of vascular features are in order. The ability to capture image quality for quality control and for informing feature computation may be crucial here. Classification of image quality using a one-dimensional image-level label, while useful, is not granular enough for some applications. Artifacts may affect only part of the image. Defocus or poor overall quality may result in a CFI usable for analysis of the main vascular arches but not any smaller vessels. A decomposition of image quality into (localized) factors may be a necessary next step in developing more robust pipelines.
In summary, our results show that our models outperform those in previous systems in every dataset and condition we evaluated, likely due to the increased size and variety of our training set. We expect that VascX models will serve as the backbone for more robust CFI vascular analysis pipelines (and other CFI-based pipelines). To serve these goals, we have made available our entire inference pipeline, including model weights and inference code with a new CFI pre-processing algorithm. With this, we aim to catalyze research towards understanding how analysis of the retinal vasculature can help us prevent, detect and monitor diseases, not only of the eye but of the rest of the body.
@@ Acknowledgments
The authors acknowledge the staff of the Eyened Reading Center for their different roles in data collection and algorithm development. Special acknowledgment to the Sinergia consortium for conceiving and supporting this project. The authors are grateful to the study participants and the staff from the Rotterdam Study, AMD-Life; the Dutch Myopia Study and to all contributors to the (publicly available) datasets that we used to train our models. This work was funded by the Swiss National Science Foundation grant no. CRSII5209510.
@@ Appendix A Dataset details
Table 4: Complementary details of the public dataset used in model development. *: 379 of 1200 samples were discarded due to containing no disc segmentation for partially visible discs. $\dagger$: 21 of 1200 CFIs were discarded due to having no associated annotation.
@@
PALM [12] 1179$\dagger$ $2124\times 2056$, $1444\times 1444$ Zeiss Visucam 500, Canon CR-2
REFUGE2 [refuge] 2000 $1634\times 1634$ $2124\times 2056$ Zeiss Visucam 500, Canon CR-2, Topcon TRC-NW400, Kowa
Rotterdam (ours) 1225 see Fig. 1
@@
REFUGE2 [refuge] 2000 $1634\times 1634$ $2124\times 2056$ Zeiss Visucam 500, Canon CR-2, Topcon TRC-NW400, Kowa
Rotterdam (ours) 10908 see Fig. 1
@@
Sample training batches
(a) Vessel segmentation.
(b) Artery-vein segmentation, where arteries, veins and crossings are separate (exclusive) classes.
Figure 6: Sample training batches of input images to vessel and artery-vein segmentation models after data augmentation, including the CFI and overlaid masks. The contrast enhanced image is not shown.
(a) Disc segmentation masks.
(b) Fovea localization heatmaps overlaid over the input images.
Figure 7: Sample training batches of input images to disc segmentation and fovea localization models after data augmentation. The contrast enhanced image is not shown.
Appendix B Feature Comparison
B.0.1 Feature implementations
In this section we describe the implementation of our vascular features, or refer to previous work for a detailed description.
Temporal Angle
We followed the implementation presented in [39] (Section 2.3 of Supplementary Material).
CRE
We followed the implementation presented in [39] (Section 2.2 of Supplementary Material). However, we measure and aggregate all vessels on the temporal side of the image that intersect the defined circle. We use only the temporal side because the nasal side is often not visible at the evaluated radii in fovea-centered images.
Vascular Density
: Defined as the fraction of 1-pixels in the binary A/V segmentation mask within the fundus mask area:
Vascular $\text{Vascular Density}=\frac{\sum_{\_}{i=1}^{H}\sum_{\_}{j=1}^{W}I(i,j)\cdot M(i,j)}{\sum_{\_}{i=1}^{H}\sum_{\_}{j=1}^{W}M(i,j)}$
where:
  • $I(i,j)$ is the binary vessel mask where each pixel is either 0 (background) or 1 (vessel).
  • $M(i,j)$ is the binary mask where the region of interest (ROI) is defined, as extracted by our pre-processing algorithm
Vessel Caliber
Vessel calibers or widths were measured per vessel segment by first fitting a spline to the segment skeleton. Perpendiculars to the spline were sampled at regular intervals. A ray casting method was used to extend the lines until they reached a background pixel (in the original artery or vein mask) on both sides, resulting in a width measurement. The median of the measurements was taken as the vessel segment caliber.
Tortuosity
Tortuosity was defined using its common definition as the length of the vessel following its points divided by its chord length (distance between its endpoints):
$T=\frac{\sum_{\_}{i=1}^{N-1}\sqrt{(x_{\_}{i+1}-x_{\_}i)^{2}+(y_{\_}{i+1}-y_{\_}i)^{2}}}{\sqrt{(x_{\_}N-x_{\_}1)^{2}+(y_{\_}N-y_{\_}1)^{2}}}$
where $(x_{\_}i,y_{\_}i)$ are the coordinates of the skeleton points, with $i=1,2,...,N$ where $N$ is the total number of points along the vessel skeleton.
Curvature
Curvature is an alternative measure of tortuosity. Given a parametrization $\gamma(t)=(x(t),y(t))$ of a twice differentiable plane curve, curvature can be expressed as:
$\kappa=\frac{|x^{\prime}y^{\prime\prime}-y^{\prime}x^{\prime\prime}|}{(x^{\prime 2}+y^{\prime 2})^{3/2}}$
where the derivatives $x^{\prime}$, $y^{\prime}$, $x^{\prime\prime}$ and $y^{\prime\prime}$ are with respect to $t$.
The mean curvature over the curve can be defined by integration of this expression:
$\bar{\kappa}=\int_{\_}0^{1}\kappa(t)$
This calculation was approximated by computing $\kappa$ at regular intervals (every 5 pixels), using splines as parametrization to approximate derivatives.
Inflection count
Another alternative measure of tortuosity, this feature counts the number of inflection points (changes of direction) in a vessel segment.
Num. Bifurcations
The number of bifurcations detected in the image. Bifurcations are detected by converting the skeleton into a graph, which is then traversed using the optic disc location as reference to defined outgoing vessels.
Bif. Angles
Smallest angle between the two child vessels in a bifurcation. A distance $\delta$ from the bifurcation point was defined, at which the angle is measured. The spline approximations are used to obtain points $(x,y)$ along the vessel segments at such $\delta$. Bifurcations with child vessels shorter than $\delta$ are ignored.
Grader $\mu$ MAE Pearson r
VascX AM LWN VascX AM LWN
Temporal Angle - A 127.49 3.44 5.40$\dagger$ 10.48$\dagger$ 0.81 0.72 0.56
Temporal Angle - V 132.46 3.02 4.85 11.68$\dagger$ 0.86 0.74 0.39
CRE - A 12.28 1.81 1.98 2.98$\dagger$ 0.68 0.38 0.22
CRE - V 19.98 1.98 2.80$\dagger$ 4.70$\dagger$ 0.74 0.58 0.14
Vasc. Density - A 4.76 0.58 1.15$\dagger$ 1.42$\dagger$ 0.77 0.61 0.48
Vasc. Density - V 5.43 0.58 1.01$\dagger$ 1.31$\dagger$ 0.79 0.63 0.39
Vessel Caliber [med] - A 5.25 1.37 1.33 1.73$\dagger$ 0.20 0.18 0.09
Vessel Caliber [std] - A 3.06 0.43 0.65$\dagger$ 0.65$\dagger$ 0.73 0.57 0.47
Vessel Caliber [med] - V 5.36 1.50 1.58 2.24$\dagger$ 0.45 0.36 0.15
Vessel Caliber [std] - V 4.81 0.49 0.89$\dagger$ 0.74$\dagger$ 0.76 0.67 0.36
Tortuosity [med] - A 1.09 0.0057 0.0064 0.008$\dagger$ 0.82 0.78 0.71
Tortuosity [med] - V 1.08 0.0044 0.0046 0.0064$\dagger$ 0.77 0.76 0.51
Curvature [med] - A 6.71 1.01 1.10 1.37$\dagger$ 0.91 0.90 0.85
Curvature [med] - V 7.83 1.19 1.31 1.34 0.81 0.77 0.76
Inflection count [med] - A 1.69 0.51 0.56 0.69$\dagger$ 0.38 0.19 0.22
Inflection count [med] - V 1.85 0.62 0.71 0.76* 0.35 0.16 0.19
Tortuosity [LW] - A 1.12 0.023 0.027$\dagger$ 0.06$\dagger$ 0.17 0.15 -0.29
Tortuosity [LW] - V 1.11 0.019 0.021$\dagger$ 0.054$\dagger$ 0.25 0.24 0.02
Bif. Angles [mean] - A 79.88 7.69 7.11 11.58$\dagger$ 0.61 0.57 0.35
Bif. Angles [med] - A 78.46 7.83 7.24* 10.75* 0.61 0.53 0.38
Bif. Angles [mean] - V 81.36 7.21 6.76 10.63$\dagger$ 0.44 0.40 0.22
Bif. Angles [med] - V 81.55 6.77 6.56 9.51$\dagger$ 0.50 0.42 0.26
Num. Bifurcations - A 15.97 6.43 6.84* 7.62$\dagger$ 0.70 0.70 0.68
Num. Bifurcations - V 18.11 7.79 7.64 9.11$\dagger$ 0.76 0.71 0.61
Table 5: Feature quality comparison on the Leuven-Haifa dataset (240 CFIs). Results of comparing MAE of features extracted using different models for artery-vein segmentation: VascX, Automorph (AM) and LittleWNet (LWN). Features extracted from ground truth segmentations were used as the reference for MAE. For vessel or bifurcation features, the aggregation function (applied on features from the whole image) is indicated in brackes: [med]: median, [std]: standard deviation, [mean]: mean value, [LW]: length-weighted; vessels were weighted by their length. The feature implementation was the same across models. Grader $\mu$ is the mean of feature values; mean MAE is the mean absolute error between ground truth features and features extracted from model outputs. Pearson coefficients are correlations between ground truth and model outputs. *: Significant difference in MAE ($p<0.05$) between the model and VascX. $\dagger$: $p<0.001$